OpenStudy (anonymous):

Simplify: ((2x+1)^(2/3)-((3(x^-1))/((2x+1)^(1/3))/((2x+1)^2)

OpenStudy (anonymous):

According to Mathematica v8, the two open parentheses in red are not balanced.

OpenStudy (anonymous):

((2x+1)^(2/3)-((3x^(-1))/(2x+1)^(1/3)))/((2x+1)^(2)) I think that's the one that's correct. I'll try to make an equation easier to read.

OpenStudy (anonymous):

It won't let me put the exponents as fractions.

OpenStudy (anonymous):

just draw it

OpenStudy (anonymous):

((2x+1)^(2/3)-((3x^(-1))/(2x+1)^(1/3)))/((2x+1)^(2))\[\left((1+2 x)^{2/3}-\left((3x{}^{\wedge}(-1))\left/(1+2 x)^{1/3}\right.\right)\right)/(1+2 x)^2 \]\[\left((1+2 x)^{2/3}-\left(\frac{3}{x}/(1+2 x)^{1/3}\right)\right)/(1+2 x)^2 \]\[\left((1+2 x)^{2/3}-\frac{3}{x (1+2 x)^{1/3}}\right)/(1+2 x)^2 \]\[\left(\frac{-3+x+2 x^2}{x (1+2 x)^{1/3}}\right)/(1+2 x)^2 \]\[\frac{-3+x+2 x^2}{x (1+2 x)^{7/3}} \]\[\frac{2 x^2+x-3}{x (2 x+1)^{7/3}} \]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Thank you for the medal.